OpenStudy (anonymous):

The Cool Company determines that the supply function for its basic air conditioning unit is S(p) = 30 + 0.006p3 and that its demand function is D(p) = 150 - 0.12p2, where p is the price. Determine the price for which the supply equals the demand.

OpenStudy (nory):

Set S(p) = D(p) and solve the equation. So 30 + 0.006p3 = 150 - 0.12p2. Solve that equation.

OpenStudy (anonymous):

@Nory i tried but it won't factor:( i dont know how to get an x without factoring it..

OpenStudy (nory):

Move all the variable terms to one side of the equation. Then multiply by 1000 to clear all the decimals. Then I think you can factor out a p2.

OpenStudy (nory):

All the constant terms have to be on the other side.

OpenStudy (anonymous):

@Nory i got 6p^3+120P^2-120000=0

OpenStudy (nory):

Move the 120000 to the other side. Then factor it. There is a common factor that you can divide both sides by.

OpenStudy (anonymous):

@Nory okay i got p=119980 and sqrt20000

OpenStudy (anonymous):

@satellite73 can you check this out??